Tia, Keith, Xiao, Cory, Backfish, Sam, Carrie, Tom
- increased intensity
- hope to keep constant intensity as MI is ramping up.
- Flux was great 5e12 running.
- Pieces machined, go out for silver plating
- experiment hall passed safety walkthrough (check some electrical labels)
- ANNIE has water circulating again, pump replaced
- DAQ being adjusted, but taking data
Near Detector Building:
- smooth running, bit of PMT noise
- 1 shifter mode in January, starting testing on this after Thanksgiving
- telephone being coordinated to receive calls from MCR
- cryo monitoring being
- NuMI open trigger running for ~1 week